OpenStudy (anonymous):

The ionization energy for potassium 419 KJ/mol. The wavelength of light emitted when an excited K atom undergoes the 4s<---- 4p transition is approximately 769 nm. Using this info, calculate the energies of the 4s and 4p orbitals in potassium.

OpenStudy (aaronq):

\(\sf \large E_{4p}= IE_{single~electron}= \dfrac{419 ~kJ/mol}{N_A}=\dfrac{419 ~kJ/mol}{6.022*10^{23}~electron/mol}\) \(\sf \large Relaxation= \Delta E_{4p\rightarrow 4s}=\dfrac{hc}{\lambda }\) \(\sf \large E_{4s}=E_{4p}+\Delta E_{4p\rightarrow 4s}\)